# Empirical analysis of Binding Precedent efficiency in the Brazilian Supreme Court via Similar Case Retrieval

### v2

@@ Abstract
Binding precedents (Súmulas Vinculantes) constitute a juridical instrument unique to the Brazilian legal system and whose objectives include the protection of the Federal Supreme Court against repetitive demands. Studies of the effectiveness of these instruments in decreasing the Court’s exposure to similar cases, however, indicate that they tend to fail in such a direction, with some of the binding precedents seemingly creating new demands. We empirically assess the legal impact of five binding precedents, 11, 14, 17, 26 and 37, at the highest court level through their effects on the legal subjects they address. This analysis is only possible through the comparison of the Court’s ruling about the precedents’ themes before they are created, which means that these decisions should be detected through techniques of Similar Case Retrieval. The contributions of this article are therefore twofold: on the mathematical side, we compare the uses of different methods of Natural Language Processing — TF-IDF, LSTM, BERT, and regex — for Similar Case Retrieval, whereas on the legal side, we contrast the inefficiency of these binding precedents with a set of hypotheses that may justify their repeated usage. We observe that the deep learning models performed significantly worse in the specific Similar Case Retrieval task and that the reasons for binding precedents to fail in responding to repetitive demand are heterogeneous and case-dependent, making it impossible to single out a specific cause.
@@ Acknowledgments
This work is the product of a collaboration between mathematicians and law specialists. As far as the latter field is concerned, in addition to the fourth author, we have received invaluable contributions from Beatriz Sabdin Chagas, Carla Marcondes Damian, Ana Clara Macedo Jaccoud and Pedro Burlini de Oliveira. As part of an introduction to research for undergraduate students, we have received support in reading documents from Livia Cales, Victoria Cury, Samanta Duarte Clara Lopes, Eduardo Portol, João Meirelles, Ana Rosenburg, and Helena Torres. Finally, we are indebted to Dr. Rafael Ramia Muneratti (Public Defender’s Office of the State of São Paulo) and Dr. Wallace Paiva Martins Junior (Public Prosecutor’s Office of the State of São Paulo), who kindly agreed to be interviewed about Binding Precedent 26. The researcher Jorge Poco is partially financed by the National Council for Scientific and Technological Development (CNPq) under Grants #311144/2022-5, the Carlos Chagas Filho Foundation for Research Support of Rio de Janeiro State (FAPERJ) under Grant #E-26/201.424/2021, and the Fundação Getulio Vargas.
@@ 1 Introduction
In recent years, the progress in Natural Language Processing (NLP) sparked significant interest in its application within the legal domain. Brazil, renowned for having the world’s highest volume of legal cases, is no exception and has already witnessed several implementations of Machine Learning methods in its judiciary system [51]. Notably, these algorithms have the potential to assist legal professionals in finding cases similar to a given one, or a given legal topic — a task commonly referred to as Similar Case Retrieval. While these methods are supported by numerous mathematical studies and proofs of concept, one could argue that some of them lack empirical validation. This article seeks to bridge this gap by comparing various document retrieval methods and evaluating their results from a legal perspective.
In particular, this article explores the legal instrument of binding precedent (Súmula Vinculante, abbreviated BP), that emerged in Brazil in the 2004 judicial reform. Its purpose was to address the issue of an overwhelming number of cases with repetitive demands inundating the Brazilian Federal Supreme Court (Supremo Tribunal Federal, STF), which ideally should handle only a limited number of cases. This situation was partly due to the Brazilian legal system’s civil law approach, where Supreme Court decisions don’t serve as authority for lower courts, leading to continued case congestion. To mitigate this, Constitutional Amendment No. 45 introduced instruments inspired by the common law system, including binding precedents. They aimed at standardizing jurisprudence, providing normative force over lower instances and the broader public administration.
After its publication, one expects the binding precedent to be cited frequently, until the legal understanding of the courts on the subject calms down, resulting in a significant decrease in the number of citations. However, among the most cited binding precedents, this trend is by no means observed. On the contrary, they show steady growth, as illustrated in Fig. 1 for the five precedents of interest in this paper: 11, 14, 17, 26 and 37. They were chosen for their high number of citations, as well as the variety of legal topics they cover, from administrative to criminal law.
The aim of this article is to shed light on the reasons why these binding precedents have not led, as expected, to a reduction in repeated demands. To this end, we employ Similar Case Retrieval methods to trace the history of these precedents and quantify certain trends. This information is subsequently used to provide a legal analysis. In summary, this article includes two main mathematical and two juridical contributions:
@@
    The application and comparative analysis of classical algorithms for Similar Case Retrieval on a database of Brazilian legal documents (including TF-IDF-based models, LSTM, BERT and regex);
@@
Documents citing a binding precedent Refer to caption
Figure 1: Histograms of the number of cases judged by the Federal Supreme Court citing Binding Precedents 11, 14, 17, 26 or 37, in our collection (Dataset #1). The bins have a length of one year, and the curves are obtained via quadratic spline interpolation. The dashed vertical lines represent the date of publication of each BP. They all exhibit an increasing trend.
@@
For our analysis, we will use a set of decisions of the Supreme Court produced between 1989 and 2018, collected and annotated by the project Supremo em Números [23]. In fact, these documents were scraped from the official STF website [^1] https://portal.stf.jus.br/, where they can be found one by one. The collection of documents gathered by the project is, however, not publicly available, and has been kindly provided to us. More precisely, we will consider two subsets of this collection, detailed further in Section 3.1: Dataset #1, consisting of all decisions citing a binding precedent (29,743 documents), and Dataset #2, gathering all decisions belonging to the topics “administrative law”, “criminal law” or “criminal procedure law” (615,262 documents). We point that the data gathered by Supremo em Números have already been used in several works [14, 13, 47]. In addition, other similar databases have been reported, such as the one maintained by the project Victor [20].
@@
The remainder of this article is organized as follows. In Section 2, we set out the legal context behind the Brazilian’s Supreme Court binding precedents, with a particular focus on the five precedents specifically studied in this article, and we give an overview of the mathematical literature surrounding Similar Case Retrieval. Datasets #1 and #2 are introduced in Section 3, as well as the models used throughout this article (TF-IDF, LSTM and BERT), that we train on the former dataset. The application of these models to the latter dataset is the focus of Section 4, where we first describe our methodology (in Section 4.1, summarized in Fig. 2), apply it to the five precedents under consideration (Sections 4.2, 4.3, 4.4, 4.5 and 4.6), and gather our findings in a juridical discussion (Section 4.7). We conclude and address future works in Section 5.
@@
Throughout this article, expressions in Portuguese will be written in italics, and may be followed by a translation in English in parenthesis. In addition, the following list contains the abbreviations that will be used, and also introduced throughout the text.
@@ 2.1 Legal context
The type of precedent known in Brazil as Súmula Vinculante (Binding Precedent) emerged in the judiciary reform of 2004, through Constitutional Amendment No. 45 (EC 45/2004, Article 103-A [^2] https://www.planalto.gov.br/ccivil_03/constituicao/Constituicao.htm#art103a), to standardize decisions, based on the importation of the jurisdiction logic usual in common law [39, p. 827]. The introduction of these instruments can be understood as an attempt to unify decision-making, aiming to achieve equality and legal certainty, i.e., to avoid that identical cases are decided in different ways, violating the constitutional guarantee of equality before the law [8, p. 20]. In fact, the new Federal Constitution Article requires, for the creation of a BP, the fulfillment of three conditions: (i) dealing with a matter under current controversy among judicial bodies, or between them and the public administration; (ii) representing a risk of serious legal uncertainty; and (iii) being the subject of significant multiplication of processes on an identical legal issue. Thus, as a legal instrument, the BP seeks not only to guarantee equality in judicial decisions, but also to increase the efficiency of the Judiciary, by avoiding delays in jurisdictional intervention.
To date, 58 BPs have been published [^3] https://portal.stf.jus.br/textos/verTexto.asp?servico=jurisprudenciaSumulaVinculante. It should be noted that, in the literature, there is a considerable criticism of the BPs, whether in terms of their institution, their formulation, or their consequences. These issues will be explored in detail throughout the article. We refer the reader to the work of [3] and [43] for a general analysis of BPs.
The precedents 11, 14, 17, 26 and 37
Once a BP is published, is it expected that the subject ceases to be a matter of controversial interpretation. Consequently, the issue should also cease to generate processes with identical demands and, primarily, prevent these from continuing to be brought before the Higher Courts. It is therefore important to establish, empirically, the degree of efficiency of this type of instrument, here understood as the degree of reduction in similar cases reaching the final stage of appeal. In this article, we analyze the impact of the creation of five binding precedents, chosen as those generating many cases, but also for the diversity of legal topics they cover. Three of them belong to criminal law (11, 14 and 26), and the others to administrative law (17 and 37). We give in the list below a brief summary of their content and refer the reader to the corresponding section for a deeper explanation, as well as the results of our analysis.
@@
    the 2008-11-12 (see Section 4.2): Determines the use of handcuffs acceptable only when some risk is anticipated, predicting disciplinary punishments to the responsible public agents and/or nullity of the penal process in the case of unjustified use.
@@
    the 2009-02-02 (see Section 4.3): Grants to the investigated individual and their attorneys full access to all documented evidence in ongoing criminal investigations.
@@
    the 2009-11-10 (see Section 4.4): In Brazil, public administration at all levels, when sentenced to pay debts resulting from juridical decisions, incorporate these amounts into the public budget by the system of precatórios (Court Order Payments), in which orders of payment are issued to beneficiaries. The BP determines that no late payment interests should be applied to the potential delay between the issuing of orders of payment and the actual payment taking place.
@@
    the 2010-02-26 (see Section 4.5): Determines the unconstitutionality of prohibiting convicts of “heinous crimes” (such as murder, rape, and drug dealing) to receive a reduction of the sentence from a fully closed to semi-open or open carceral system. Moreover, the BP authorizes judges to employ criminological exams to determine the feasibility of this reduction of the sentence.
@@
    the 2014-10-23, (see Section 4.6): Establishes that the Judiciary cannot increase salaries of public servants based on the principle of isonomy (equality) without prior legislation to authorize such adjustments.
@@ 2.2 NLP for legal documents
The literature surrounding NLP for legal document analysis is rich, reflecting the growing interest in leveraging computational techniques to navigate the complexities of legal texts. Well-studied topics in this area include the analysis of the precedent network [36, 14, 41], Named Entity Recognition [13], summarization of legal texts [28, 26] and prediction of judicial outcome [2, 4, 5]. In addition, of particular interest to us is the detection of similar documents, reviewed further in the next paragraph.
@@
To tackle these problems, a standard technique consists of embedding (i.e., vectorizing) the documents, the most classical methods being bag-of-words, TF-IDF, and $n$-grams. Among recent techniques of words and documents embeddings, we can cite GLoVE [42], contextualized word representation [44], word2vec [12], doc2vec [33], Latent Dirichlet Allocation (LDA) [7], entities and relations-based embedding [72], Universal Sentence Encoder (USE) [10] and TextCNN [11]. Another alternative are the pre-trained language models, such as BERT [21] and its variants [75, 45, 29, 67].
@@
In particular, word2vec is used by [24] to study appellate court modifications in Brazilian legal documents, that is, modifcations by the Supreme Court of the lower Court judge’s decision. On the other hand, TF-IDF has been already used for the analysis of binding precedents, and has been reported to outperform other embeddings (such as doc2vec and USE) [47]. We will consider, in this article, TF-IDF embeddings coupled with several classifiers, as well as a recurrent neural network (LSTM) and a Large Language Model (BERT).
Automatic Similar Case Matching and Retrieval
@@
At their most elementary level, although having a specific binding property, BPs are precedent and, as such, the task of searching for documents similar to the object of a BP, in content and language, may be modeled through textual proximity and/or classification techniques. The problem of identifying juridical decisions similar in a corpus of documents, known as automatic Similar Case Matching, has received significant attention in the literature. For example, a multitude of NLP embedding models have been compared in the CAIL2019-SCM dataset, a corpus of thousands of decisions published by the Supreme People’s Court of China, in which the task is to determine, for a triple of cases $(A,B,C)$, whether $A$ is more similar to $B$ or to $C$ [71]. Because of the problem structure, different NLP techniques have been applied to solving the task, including pre-trained models [71], context-based multi-learning [19], ideas from optimal transport [73], casual inference [59], and even techniques combining regex and neural networks [30]. Further techniques for Similar Case Matching were developed and applied to datasets other than CAIL [32, 38], with the work of [6] especially significant for mixing both NLP and citation-based approaches.
@@
Although related in spirit, Similar Case Matching is not exactly the task we are tackling. Instead of computing similarity indexes between pairs of documents or telling which of two documents is more similar to a third one, we need to find in a huge unlabeled corpus of decisions all decisions that are similar in content to a particular document. This alternative problem, arguably harder, is known as Similar Case Retrieval and unsurprisingly, has been tackled through different artificial intelligence models [58, 46, 68, 65], including some that we will be using in this work [47]. Surprisingly, though, none of these works seemed to use these techniques to describe the temporal behavior of specific precedents as we do here, restraining themselves to tools for aiding the work of juridical agents.
@@
The reason why we claim the task of case retrieval to be harder than case matching goes beyond the natural computational problem of the sheer size of inputs in the former. Indeed, in this context, when going from the training set (called Dataset #1 in this article), containing actual citations of BPs, to the test set (Dataset #2), where no citations to BPs are to be found, some overfitting of the trained models is expected. However, this bias cannot be directly measured, as there are no documents of this second set which are labeled. Notice that the difficulty is twofold: not only there will be different proportions of positive and negative classes between the two sets (the positive class representing documents citing a specific BP), but also the very textual elements from the positive and negative classes are expected to be distinct, even when regex is used to remove the citations themselves. In conclusion, the problem of Similar Case Retrieval tackled in this article goes beyond other analyses of classification problems found in the literature, where one can, a posteriori, quantify the validity of models through accuracy, recall, or $F_{1}$ score. In order to validate our results, we will propose a thorough legal analysis, based in particular on the reading of documents.
@@
Although there is a multitude of NLP models that solve both the similarity matching and retrieval problems, none of these have yet been applied to the actual juridical study of the creation, the evolution, and the authority of precedents. On the other hand, empirical legal literature has often investigated these questions using ideas from graph theory. That is, given the natural referential structure inherited from the use of precedents, a substantial part of the empirical study of courts is dedicated to modeling them through directed graphs, generated by the set of all decisions, and where there is an arrow from a decision $x$ towards a decision $y$ if and only if $y$ cites $x$ as a precedent. Drawing from the general theory of citation networks [31], authors have developed metrics to measure the authority of a specific decision as a precedent in a determined court through the decision’s (outward) degree in the induced graph, emulating the hypothesis that the more important a precedent, the more cited it is.
These metrics have then been applied to identify the most important precedents in different settings and the evolution of their authority through time. In a fundamental empirical work on the United States Supreme Court [25], for example, the authors used the (normalized) degree of the decisions to identify historical features of the most important precedents to the court, concluding that for most cases, when an important decision is taken by the court, the number of documents citing it follows a curve which tends to increase with time, reach a peak, and then decrease exponentially. Interestingly, the authors indicate that although this seems to be a general trend, internal forces within the court, such as the reversion of previous decisions, may significantly impact these curves. The authors of [69] point that even external sources, such as the court’s political composition impact the citation curves of precedents in the Supreme Court and, in a study considering all instances in the American justice, [55] identifies the probability distributions generating the precedent graphs. Similar conclusions regarding the change of authority over time are found about the International Criminal Court [60], the Canadian legal system [40], and the Court of Justice of the European Union [61]. In this last study, the authors had access to a list of the most important precedents identified by the court’s legal experts and attested that, in practice, there exist important cases, called symbolic, which will not have high authority scores, even considering their doctrinal impact.
@@
Network analysis of Brazilian system

### v3

# Empirical analysis of binding precedent efficiency in Brazilian Supreme Court via case classification

@@ Abstract
Binding precedents (súmulas vinculantes) constitute a juridical instrument unique to the Brazilian legal system and whose objectives include the protection of the Federal Supreme Court against repetitive demands. Studies of the effectiveness of these instruments in decreasing the Court’s exposure to similar cases, however, indicate that they tend to fail in such a direction, with some of the binding precedents seemingly creating new demands. We empirically assess the legal impact of five binding precedents, 11, 14, 17, 26, and 37, at the highest Court level through their effects on the legal subjects they address. This analysis is only possible through the comparison of the Court’s ruling about the precedents’ themes before they are created, which means that these decisions should be detected through techniques of Similar Case Retrieval, which we tackle from the angle of Case Classification. The contributions of this article are therefore twofold: on the mathematical side, we compare the use of different methods of Natural Language Processing — TF-IDF, LSTM, Longformer, and regex — for Case Classification, whereas on the legal side, we contrast the inefficiency of these binding precedents with a set of hypotheses that may justify their repeated usage. We observe that the TF-IDF models performed slightly better than LSTM and Longformer when compared through common metrics; however, the deep learning models were able to detect certain important legal events that TF-IDF missed. On the legal side, we argue that the reasons for binding precedents to fail in responding to repetitive demand are heterogeneous and case-dependent, making it impossible to single out a specific cause. We identify five main hypotheses, which are found in different combinations in each of the precedents studied.
@@ Acknowledgments
This work is the product of a collaboration between mathematicians and Law specialists. As far as the latter field is concerned, in addition to the fourth author, we have received invaluable contributions from Beatriz Sabdin Chagas, Carla Marcondes Damian, Ana Clara Macedo Jaccoud, and Pedro Burlini de Oliveira. As part of an introduction to research for undergraduate students, we have received support in reading documents from Livia Cales, Victoria Cury, Samanta Duarte Clara Lopes, Eduardo Portol, João Meirelles, Ana Rosenburg, and Helena Torres. We are also particularly grateful to José Luiz Nunes for his help in accessing and navigating the dataset, as well as his sound methodological recommendations. Finally, we are indebted to Dr. Rafael Ramia Muneratti (Public Defender’s Office of the State of São Paulo) and Dr. Wallace Paiva Martins Junior (Public Prosecutor’s Office of the State of São Paulo), who kindly agreed to be interviewed about Binding Precedent 26. The researcher Jorge Poco is partially financed by the National Council for Scientific and Technological Development (CNPq) under Grants #311144/2022-5, the Carlos Chagas Filho Foundation for Research Support of Rio de Janeiro State (FAPERJ) under Grant #E-26/201.424/2021, and the Fundação Getulio Vargas. The researcher Jean R. Ponciano is partially financed by the University of São Paulo (PRPI Ordinance No. 1032, Process 22.1.09345.01.2). Lucas Resck acknowledges financial support from Fundação Getulio Vargas and partial support from the Coordination for the Improvement of Higher Education Personnel (CAPES).
@@ 1 Introduction
In recent years, the progress in Natural Language Processing (NLP) sparked significant interest in its application within the legal domain. Brazil, renowned for having the world’s highest volume of legal cases, is no exception and has already witnessed several implementations of Machine Learning methods in its judiciary system [24, 35, 44, 15, 14, 25, 5, 48, 42, 6, 52]. Notably, these algorithms have the potential to assist legal professionals in finding cases similar to a given one, or a given legal topic — a task commonly referred to as Similar Case Retrieval. While these methods are supported by numerous mathematical studies and proofs of concept, one could argue that some of them lack empirical validation. This article seeks to bridge this gap by comparing various document retrieval methods and evaluating their results from a legal perspective.
In particular, we explore the legal instrument of binding precedent (súmula vinculante, abbreviated BP), that emerged in Brazil in the 2004 judicial reform. Its purpose was to address the issue of an overwhelming number of cases with repetitive demands inundating the Brazilian Federal Supreme Court (Supremo Tribunal Federal, STF), which ideally should handle only a limited number of cases. This situation was partly due to the Brazilian legal system’s civil law approach, where Supreme Court decisions do not serve as authority for lower courts, leading to continued case congestion. To mitigate this, Constitutional Amendment 45 introduced instruments inspired by the common law system, including binding precedents. They aimed at standardizing jurisprudence, providing normative force over lower instances and the broader public administration.
After its publication, one expects the binding precedent to be cited frequently, until the legal understanding of the courts on the subject settles down, resulting in a significant decrease in the number of citations. However, among the most cited binding precedents, this trend is by no means observed. On the contrary, they show steady growth, as illustrated in Fig. 1 for the five precedents of interest in this paper: 11, 14, 17, 26, and 37. They were chosen for their high number of citations, as well as the variety of legal topics they cover, from administrative to criminal law.
Number of documents citing a binding precedent Refer to caption
Figure 1: Histograms of the number of cases judged by the Federal Supreme Court citing Binding Precedents 11, 14, 17, 26, or 37, in our collection (Dataset #1). The bins have a length of one year, and the curves are obtained via quadratic spline interpolation. The dashed vertical lines represent the date of publication of each BP. We draw the reader’s attention to the fact that they all exhibit an increasing trend.
This article aims to shed light on the reasons why these precedents have not led, as expected, to a reduction in repeated demands. To this end, we employ Similar Case Retrieval and Case Classification methods to trace the history of these precedents and quantify certain trends. This information is subsequently used to provide a legal analysis. In summary, this article includes two main mathematical and two juridical contributions:
@@
    The application and comparative analysis of classical algorithms for Case Classification on a database of Brazilian legal documents (including TF-IDF-based models, LSTM, Longformer, and regex);
@@
For our analysis, we used a set of decisions of the Supreme Court produced between 1989 and 2018, collected and annotated by the project Supremo em Números [24]. In fact, these documents were scraped from the official STF website, where they can be found one by one. The collection of documents gathered by the project is, however, not publicly available, and has been kindly provided to us. More precisely, we will consider two subsets of this collection, detailed further in Section 3.1: Dataset #1, consisting of all decisions citing a binding precedent (29,743 documents), and Dataset #2, gathering all decisions belonging to the topics “Administrative Law”, “Criminal Law” or “Criminal Procedure Law” (634,068 documents). We point out that the data gathered by Supremo em Números have already been used in several works [15, 14, 48]. In addition, other similar databases have been reported, such as the one maintained by the project Victor [21].
@@
The remainder of this article is organized as follows. In Section 2, we set out the legal context behind the Brazilian Supreme Court’s binding precedents, with a particular focus on the five precedents specifically studied in this article, and we give an overview of the mathematical literature surrounding Similar Case Retrieval. Datasets #1 and #2 are introduced in Section 3, as well as the models used throughout this article (TF-IDF, LSTM, and Longformer), that we train and test on the former dataset. In Section 4, we apply these models to the latter dataset and evaluate the quality of their predictions. The legal analysis of our results is the focus of Section 5, where we first describe our methodology (in Section 5.1, summarized in Fig. 2 below), apply it to the five precedents under consideration (Sections 5.2, 5.3, 5.4, 5.5 and 5.6), and gather our findings in a juridical discussion (Section 5.7). We conclude and address future works in Section 6.
@@
Throughout this article, expressions in Portuguese will be written in italics and may be followed by a translation in English in parenthesis. In addition, the following list contains the abbreviations that will be used, and also introduced throughout the text.
@@ 2.1 Legal context
The type of precedent known in Brazil as súmula vinculante (binding precedent) emerged in the judiciary reform of 2004, through Constitutional Amendment 45 (EC 45/2004, Article 103-A [^1] Article 103-A https://www.planalto.gov.br/ccivil_03/constituicao/Constituicao.htm#art103a), to standardize decisions, based on the importation of the jurisdiction logic usual in common law [40, p. 827]. The introduction of these instruments can be understood as an attempt to unify decision-making, aiming to achieve equality and legal certainty, i.e., to avoid identical cases being decided in different ways, violating the constitutional guarantee of equality before the law [9, p. 20]. In fact, the new Federal Constitution article requires, for the creation of a BP, the fulfillment of three conditions: (i) dealing with a matter under current controversy among judicial bodies or between them and the public administration; (ii) representing a risk of serious legal uncertainty; and (iii) being the subject of significant multiplication of processes on an identical legal issue. Thus, as a legal instrument, the BP seeks not only to guarantee equality in judicial decisions but also to increase the efficiency of the Judiciary, by avoiding delays in jurisdictional intervention.
To date, 58 BPs have been published [^2] List of BPs https://portal.stf.jus.br/textos/verTexto.asp?servico=jurisprudenciaSumulaVinculante. It should be noted that, in the literature, there is considerable criticism of the BPs, whether in terms of their institution, their formulation, or their consequences. These issues will be explored in detail throughout the article. We refer the reader to the work of [3] and [44] for a general analysis of BPs.
The precedents 11, 14, 17, 26, and 37
Once a BP is published, it is expected that the subject ceases to be a matter of controversial interpretation. Consequently, the issue should also cease to generate processes with identical demands and, primarily, prevent these from continuing to be brought before the Higher Courts. It is therefore important to establish, empirically, the degree of efficiency of this type of instrument, understood here as the degree of reduction in similar cases reaching the final stage of appeal. In this article, we analyze the impact of the creation of five binding precedents, chosen as those generating many cases, but also for the diversity of legal topics they cover. Three of them belong to criminal law (11, 14, and 26), and the others to administrative law (17 and 37). We give in the list below a brief summary of their content and refer the reader to the corresponding section for a deeper explanation, as well as the results of our analysis.
@@
    2008-11-12 (see Section 5.2): Determines the use of handcuffs acceptable only when some risk is anticipated, predicting disciplinary punishments to the responsible public agents and/or nullity of the penal process in the case of unjustified use.
@@
    2009-02-02 (see Section 5.3): Grants to the investigated individual and their attorneys full access to all documented evidence in ongoing criminal investigations.
@@
    2009-11-10 (see Section 5.4): In Brazil, public administration at all levels, when sentenced to pay debts resulting from juridical decisions, incorporate these amounts into the public budget by the system of precatórios (court order payments), in which orders of payment are issued to beneficiaries. The BP determines that no late payment interests should be applied to the potential delay between the issuing of orders of payment and the actual payment taking place.
@@
    2010-02-26 (see Section 5.5): Determines the unconstitutionality of prohibiting convicts of “heinous crimes” (such as murder, rape, and drug dealing) to receive a reduction of the sentence from a fully closed to semi-open or open carceral system. Moreover, the BP authorizes judges to employ “criminological examinations” to determine the feasibility of this reduction of the sentence.
@@
    2014-10-23 (see Section 5.6): Establishes that the Judiciary cannot increase salaries of public servants based on the principle of isonomy (equality) without prior legislation to authorize such adjustments.
@@ 2.2 NLP for legal documents
The literature surrounding NLP for legal document analysis is rich, reflecting the growing interest in leveraging computational techniques to navigate the complexities of legal texts. Well-studied topics in this area include the analysis of the precedent network [37, 15, 42], Named Entity Recognition [14], summarization of legal texts [29, 27], and prediction of judicial outcome [2, 5, 6]. In addition, of particular interest to us is the detection of similar documents, reviewed further in the next paragraph.
@@
To tackle these problems, a standard technique consists of embedding (i.e., vectorizing) the documents, the most classical methods being bag-of-words, TF-IDF, and $n$-grams. Among recent techniques of words and documents embeddings, we can cite GLoVE [43], contextualized word representation [45], word2vec [13], doc2vec [34], Latent Dirichlet Allocation (LDA) [8], entities and relations-based embedding [75], Universal Sentence Encoder (USE) [11], and TextCNN [12]. Another alternative is the pre-trained language models, such as BERT [22] and its variants [77, 46, 30, 4, 70].
@@
In particular, word2vec is used by [25] to study appellate court modifications in Brazilian legal documents, that is, modifications by the Supreme Court of the lower Court judge’s decision. On the other hand, TF-IDF has been already used for the classification of documents citing binding precedents and has been reported to outperform other embeddings (such as doc2vec, USE, and Longformer) [48]. In this article, we will consider TF-IDF embeddings coupled with several classifiers, as well as a recurrent neural network (LSTM) and a Large Language Model (Longformer).
Similar Case Matching, Retrieval, and Case Classification
@@
At their most elementary level, although having a specific binding property, BPs are precedent and, as such, the task of searching for documents similar to the object of a BP, in content and language, may be modeled through textual proximity and/or classification techniques. The problem of identifying juridical decisions similar in a corpus of documents, known as automatic Similar Case Matching, has received significant attention in the literature. For example, a multitude of NLP embedding models have been compared in the CAIL2019-SCM dataset, a corpus of thousands of decisions published by the Supreme People’s Court of China, in which the task is to determine, for a triple of cases $(A,B,C)$, whether $A$ is more similar to $B$ or to $C$ [74]. Because of the problem structure, different NLP techniques have been applied to solving the task, including pre-trained models [74], context-based multi-learning [20], ideas from optimal transport [76], causal inference [61], and even techniques combining regex and neural networks [31]. Further techniques for Similar Case Matching were developed and applied to datasets other than CAIL [33, 39], with the work of [7] especially significant for mixing both NLP and citation-based approaches.
@@
Although related in spirit, Similar Case Matching is not exactly the task we are tackling. Instead of computing similarity indexes between pairs of documents or telling which of two documents is more similar to a third one, we need to find in a huge unlabeled corpus of decisions all decisions that are similar in content to a particular document. This alternative problem, arguably harder, is known as Similar Case Retrieval and unsurprisingly, has been tackled through different artificial intelligence models [60, 47, 71, 68], including some that we will be using in this work [48]. Curiously, though, none of these works seemed to use these techniques to describe the temporal behavior of specific precedents as we do here, restraining themselves to tools for aiding the work of juridical agents.
In order to retrieve documents similar to the thematic content of a binding precedent, we will employ a trick that reduces the problem to that of Case Classification. Starting from the collection of all documents (Dataset #2), we will consider the subset of those citing a BP (Dataset #1). Based on this second dataset, which is labeled (each document is associated with a BP), we can simply train models for this classification task, using any of the various methods popular in the literature. Subsequently, we will apply these models to the larger dataset, thus addressing the problem of Similar Case Retrieval. In short, our approach is hybrid, and we will use the term “Case Classification” to describe it in the remainder of this article.
@@
To close this section, we emphasize that the reason why we claim the task of case retrieval to be harder than case matching goes beyond the natural computational problem of the sheer size of inputs in the former. Indeed, in this context, when going from the training set (Dataset #1), containing actual citations of BPs, to the complete set (Dataset #2), where no citations to BPs are to be found, some overfitting/underfitting of the trained models is expected. However, this bias cannot be directly measured, as the documents of this second dataset are not labeled. Notice that the difficulty is twofold: not only there will be different proportions of positive and negative classes between the two sets (the positive class representing documents citing a specific BP), but also the very textual elements from the positive and negative classes are expected to be distinct, even when regex is used to remove the citations themselves. In conclusion, the problem of Similar Case Retrieval tackled in this article goes beyond other analyses of classification problems found in the literature, where one can, a posteriori, quantify the validity of models through accuracy, recall, or $F_{1}$ score. To validate our results, two techniques will be used. First, we will employ the labels in Dataset #1 (which accounts for 4.7% of Dataset #2) to compute common metrics such as those cited above. This will be performed in Section 3. Secondly, we will propose a thorough manual validation, based in particular on the reading of documents and the analysis of prediction curves (in Section 4).
@@
Although there is a multitude of NLP models that solve both the similarity matching and retrieval problems, to our knowledge, none of these have yet been applied to the actual juridical study of the creation, the evolution, and the authority of precedents. On the other hand, empirical legal literature has often investigated these questions using ideas from graph theory. That is, given the natural referential structure inherited from the use of precedents, a substantial part of the empirical study of courts is dedicated to modeling them through directed graphs, with vertices the set of all decisions, and with an arrow from a decision $x$ towards a decision $y$ if and only if $y$ cites $x$ as a precedent. Drawing from the general theory of citation networks [32], authors have developed metrics to measure the authority of a specific decision as a precedent in a determined court through the decision’s (outward) degree in the induced graph, emulating the hypothesis that the more important a precedent, the more cited it is.
These metrics have then been applied to identify the most important precedents in different settings and the evolution of their authority through time. In a fundamental empirical work on the United States Supreme Court [26], for example, the authors used the (normalized) degree of the decisions to identify historical features of the most important precedents to the court, concluding that for most cases, when an important decision is taken by the court, the number of documents citing it follows a curve which tends to increase with time, reach a peak, and then decrease exponentially. Interestingly, the authors indicate that although this seems to be a general trend, internal forces within the court, such as the reversion of previous decisions, may significantly impact these curves. The author of [72] points out that even external sources, such as the court’s political composition impact the citation curves of precedents in the Supreme Court and, in a study considering all instances in the American justice, [57] identifies the probability distributions generating the precedent graphs. Similar conclusions regarding the change of authority over time are found in the International Criminal Court [63], the Canadian legal system [41], and the Court of Justice of the European Union [64]. In this last study, the authors had access to a list of the most important precedents identified by the court’s legal experts and attested that, in practice, there exist important cases, called symbolic, which will not have high authority scores, even considering their doctrinal impact.
@@
Network analysis of the Brazilian system